What are the corresponding degrees of freedom and critical value for a one-way ANOVA with 4 conditions, 6 subjects in each condition, and α = 0.05?

BG:  
WG:  
Total:  
critical value:
Answer #1

Answer:

Given,

Conditions = k = 4

Subjects = 6

n1 = n2 = n3 = n4 = 6

No. of observations = n = 24

Now,

Between groups = k - 1

= 4 - 1

BW = 3

Within groups = n - k

= 24 - 4

WG = 20

Total = n - 1

= 24 - 1

Total = 23

Now critical value for a one way ANOVA is

F 0.05,3,20 = 3.0984